OpenStudy (anonymous):

Find (partial derivatives) dz/dv, if z= x^2 + xy^3 where x = uv^2 + w^3 and y = u + ve^w, u = 2, v=1, w=0

ganeshie8 (ganeshie8):

\(\large \frac{\partial z}{\partial v} = \frac{\partial z}{\partial x} \frac{\partial x}{\partial v} +\frac{\partial z}{\partial y} \frac{\partial y}{\partial v} \)

ganeshie8 (ganeshie8):

find partials of z, with respect to x and y find partials fo x and y with respect to v plugin

OpenStudy (anonymous):

can you show me working out plz?

ganeshie8 (ganeshie8):

\(\large z = x^2 + xy^3\) \(\large \frac{\partial z}{\partial x} = 2x + y^3\) \(\large \frac{\partial z}{\partial y} = 3xy^2\)

ganeshie8 (ganeshie8):

\(\large x = uv^2 + w^3 \) \(\large y = u + ve^w \) \(\large \frac{\partial x}{\partial v} = 2uv\) \(\large \frac{\partial y}{\partial v} = e^w\)

ganeshie8 (ganeshie8):

plug them

OpenStudy (mathmale):

@bizpro: I found it easier to write out a solution on paper. See the attached image. Please ask all the questions you want. MM

OpenStudy (anonymous):

thnx :)